A Bibliography of Publications of Hartmut Surmann

Hartmut Surmann

GMD, German National Research Center of Computer Science

SET Schloss Birlinghoven 53754 Sankt Augustin

Germany

Tel: +49 2241 14-2518 FAX: +49 2241 14-2598

E-mail: surmann@gmd.de (Internet)

21 November 2023 Version 1.09

Abstract

This bibliography records publications of Hartmut Surmann.

Title word cross-reference

3 [NLHS07]. 6 [NLHS07].

Akkumulatoren [Sur92, SF92]. Algorithms [SKG93]. Algorithmus [KSG94]. analoger [GHK^{+}93]. analogue [LSHP95]. analogue/digital [LSHP95]. approaches [Sur01]. Architecture [STUG92, SUG93]. Architektur [SUG92]. Architekturkonzept [URSG92]. Arrays [STUG92, SUG93]. associative [RS91]. assoziativer [Sur89]. attentive [FRNS03]. ausgewählter [Sur89]. Automatic [SKG93]. Automatischer [Sur95]. based [STUG92, SMG92, SUKG94, SU95]. basierend [Sur89]. Basis [Sur92]. Batterien [FOWS94]. batteries [SF92]. Baustein [GHK^{+}93]. binary [RS91].

charging [SF92]. Chrom [Sur94]. Circuits [SKRG91]. Clevere [GS92]. Cobalt [Sur94]. Cobalt-Chrom-Dentalgußlegierungen [Sur94]. combination [Sur01]. Control [SKG93]. Controller [STUG92, SHH^{+}92, SUG93, SF92, SHP95, GRS^{+}91, SF92, SUG92]. Controllers [Sur92]. controlling [SHP95].

D [NLHS07, NLHS07]. Datenraten [SUG92]. Decision [SKG93]. Defect [SKRG91]. Dentalgußlegierungen [Sur94]. Design [SKG93]. digital [LSHP95]. digitaler [GHK^{+}93]. distributed [SMG92]. driven [SF92]. Durchführung [Sur94].

Entwicklungsumgebung [SHH^{+}92]. entworfen [GS92]. Entwurf [GHK^{+}93, Sur95]. environments [NLHS07]. eye [FRNS03].

faults [RS91]. Fehlertoleranz [Sur89]. Field [STUG92, SUG93]. First [SKRG91]. functional [Sur01]. Fuzzy [GHK^{+}93, HSP95, KSG94, STUG92, SHH^{+}92, SUG93, SKG93, Sur95, SU95, URSG92, LSHP95, SMG92, SUKG94, SHP95, Sur01, GRS^{+}91, RHK^{+}92, SF92, SUG92, Sur92]. Fuzzy-Controller [STUG92, SHH^{+}92, GRS^{+}91, SF92, SUG92]. Fuzzy-Controllers [Sur92]. Fuzzy-RISC-Prozessors [URSG92]. Fuzzy-rule-based [SU95]. Fuzzy-Systeme [GHK^{+}93, RHK^{+}92]. Fuzzy-Systemen [KSG94].

Gate [STUG92, SUG93]. general [SU95]. Genetic [SKG93]. genetischer [KSG94]. German [SF92]. gesteuertes [SF92]. Glänzen [Sur94]. Glänzgerät [Sur94]. Guiding [SHP95].

Hardware [Sur89, SUKG94, GRS^{+}91]. Hardware-Fehlertoleranz [Sur89]. hohe [SUG92].

indoor [LSHP95]. Integrated [SKRG91]. Integrationsgerechter [GHK^{+}93].

kind [SUKG94]. Komponente [SHH^{+}92]. künstlichen [RHK^{+}92].

Ladeverfahren [SF92, Sur92]. laser [FRNS03].

mapping [NLHS07]. Masks [SKRG91]. memory [RS91]. Mikroelektronik [RHK^{+}92]. mixed [LSHP95]. Mobile [HSP95, LSHP95, SHP95]. modal [FRNS03]. modes [SF92]. multi [FRNS03]. multi-modal [FRNS03].

Navigation [HSP95, LSHP95]. necessary [SUKG94]. Nets [SKRG91]. Netze [RHK^{+}92]. Netzen [Sur89]. Neural [SKRG91]. neuronalen [RHK^{+}92, Sur89]. neuronaler [SHH^{+}92]. NiCd [Sur92, FOWS94, SF92]. NiCd-Akkumulatoren [Sur92, SF92]. NiCd-Batterien [FOWS94]. Nochmals [KSG94].

Optimieren [KSG94]. optimiertem [URSG92]. Optimized [SUG93]. Organizing [SKG93, SMG92]. outdoor [NLHS07].

Pipeline [SUG92]. processors [SU95]. Programmable [STUG92, SUG93]. Prozessors [URSG92]. purpose [SU95].

quick [SF92].

Realisierung [GHK^{+}93]. Realtime [HSP95]. reasoning [Sur01]. Recognition [SKRG91]. regelbasierten [KSG94]. Regler [GS92]. Results [SKRG91]. RISC [URSG92]. robot [LSHP95]. Robots [HSP95, SHP95]. rule [SMG92, SUKG94, SU95]. rule-based [SMG92].

Schnell [SF92, GS92, Sur92]. Schnell-Ladeverfahren [SF92, Sur92]. Schnelladeverfahren [FOWS94]. Self [SKG93, SMG92]. Self-Organizing [SKG93, SMG92]. Simulation [GHK^{+}93]. SLAM [NLHS07]. sowie [Sur94]. Speicherbedarf [URSG92]. Speicherkonzepte [Sur89]. stuck [RS91]. stuck-at-faults [RS91]. System [HSP95, LSHP95, SMG92, SUKG94]. Systeme [GHK^{+}93, RHK^{+}92]. Systemen [KSG94, Sur95]. Systems [SKG93, SU95].

Tolerance [RS91].

Untersuchung [Sur89].

Verfahren [Sur94]. Verfahrens [Sur94]. VLSI [SUG92]. VLSI-Pipeline [SUG92].

zum [GHK^{+}93, KSG94, Sur94]. zur [Sur94].

References

Flinspach:1994:SFN

[FOWS94] G. Flinspach, A. Osswald, P. Wolf, and H. Surmann. Schnelladeverfahren für NiCd-Batterien, volume 2. Anwendungen, chapter 9. Oldenbourg-Verlag, München, Wien, Oldenbourg, 1994. ISBN 3-486-22677-0. 2. Fuzzy-Praxistagung, Augsburg, 16./17. März 1993.

Frintrop:2003:AMM

[FRNS03] Simone Frintrop, Erich Rome, Andreas Nüchter, and Hartmut Surmann. An attentive, multi-modal laser “eye”. In Computer vision systems. Third international conference, ICVS 2003, Graz, Austria, April 1–3, 2003. Proceedings, pages 202–211. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 2003. ISBN 3-540-00921-3 (paperback).

Goser:1993:IER

[GHK^{+}93] K. Goser, Ch. Heite, A. Kanstein, T. Kettner, K. Schumacher, H. Surmann, and A. Ungering. Integrationsgerechter Entwurf und Realisierung analoger und digitaler Fuzzy-Systeme: Von der Simulation zum Baustein. Universität Dortmund, Fakultät für Elektrotechnik, Dortmund, Germany, August 24, 1993. ISSN 0941-4169.

Goser:1991:HFF

[GRS^{+}91] K. Goser, U. Rückert, K. Schumacher, A. Ungering, C. Heite, and Hartmut Surmann. Hardware für Fuzzy-Controller. In Proceedings 1. Dortmund Fuzzy Tage, Dortmund, 1991.

Goser:1992:CRS

[GS92] Karl Goser and Hartmut Surmann. Clevere Regler schnell entworfen. Elektronik, pages 60–68, July 1992.

Huser:1995:FSR

[HSP95] Jörg Huser, Hartmut Surmann, and Liliane Peters. A fuzzy system for realtime navigation of mobile robots. In Information Processing for Intelligent Service Robots, Workshop 8, KI-95, Bielefeld, pages 170–172, 11–12 September

1995. URL http://borneo.gmd.
de/EIA/papers/robot_papers.
html.

Kanstein:1994:NGA

[KSG94] Andreas Kanstein, Hartmut Surmann, and Karl Goser. Nochmals ein genetischer Algorithmus zum Optimieren von regelbasierten Fuzzy-Systemen? In B. Reusch, editor, Fuzzy Logik—Theorie und Praxis, pages 86–93. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1994.

Lund:1995:MAD

[LSHP95] Trevor Lund, Hartmut Surmann, Jörg Huser, and Liliane Peters.

A mixed analogue/digital fuzzy system for indoor mobile robot navigation. In Third Australian an New Zealand Conference on Intelligent Information Systems, ANZIIS 95, Vol. 1, pages 204–209, 27 November–1 December 1995.

Nuchter:2007:SMO

[NLHS07] Andreas Nüchter, Kai Lingemann, Joachim Hertzberg, and Hartmut Surmann. 6D SLAM — 3D mapping outdoor environments. Journal of Field Robotics, 24(8–9):699– 722, 2007. ISSN 1556-4959 (print), 1556-4967 (electronic).

Rueckert:1992:MKN

[RHK^{+}92] U. Rückert, D. Heimann, I. Kreuzer, M. Mostardt-Lohbeck, Ch. Pintaske, S. Rüping, H. Surmann, V. Tryba, and Ansgar Ungering. Mikroelektronik der künstlichen neuronalen Netze und der Fuzzy-Systeme. In 20 Jahre Fakultät für Elektrotechnik, Fakultät für Elektrotechnik, pages 71–79, 1992. ISSN 0941-4169.

Rueckert:1991:TBA

[RS91] Ulrich Rückert and Hartmut Surmann. Tolerance of a binary associative memory towards stuckat-faults. In Proceedings Int. Conf. on Artificial Neural Networks, ICANN-91 Espoo, Finland, pages 1195–1198, 24–28 July 1991.

Surmann:1992:FCG

[SF92] Hartmut Surmann and Gunter Flinspach. Fuzzy-Controller gesteuertes Schnell-Ladeverfahren für NiCd-Akkumulatoren. (German)

[Fuzzy controller driven with quick charging modes for NiCd batteries]. VDE-Fachtagung: “Technische Anwendungen von Fuzzy-Systemen”, Dortmund, pages 159– 168, 12–13 November 1992.

Surmann:1992:EFF

[SHH^{+}92] Hartmut Surmann, Kai Heesche, Martin Hoh, Karl Goser, and Rainer Rudolf. Entwicklungsumgebung für Fuzzy-Controller mit neuronaler Komponente. VDE-Fachtagung: “Technische Anwendungen von Fuzzy-Systemen”, Dortmund, pages 288–297, 12– 13 November 1992.

Surmann:1995:GCM

[SHP95] Hartmut Surmann, Jörg Huser, and Liliane Peters. Guiding and controlling mobile robots with a fuzzy controller. In Fourth IEEE International Conference on Fuzzy Systems, Yokohama, Japan, pages 83–88, 20–24 March

1995. URL http://borneo.gmd.
de/EIA/papers/robot_papers.

html. Distinguished with Robot Intelligence Award.

Surmann:1993:SOG

[SKG93] Hartmut Surmann, Andreas Kanstein, and Karl Goser. Self-organizing and genetic algorithms for an automatic design of fuzzy control and decision systems. In Proceedings of the First European Congress on Fuzzy and Intelligent Technologies, EUFIT’93, Aachen, pages 1097–1104, 7–10 September

1993. URL http://borneo.gmd.
de/EIA/papers/robot_papers.
html.

Surmann:1991:NND

[SKRG91] Hartmut Surmann, Benhür Kiziloglu, Ulrich Rückert, and Karl Goser. Neural nets for defect recognition on masks and integrated circuits: First results. In Neural Networks & their Applications: Proceedings of Neuro-Nimes ’91, pages 581– 591, 1991.

Surmann:1992:DSO

[SMG92] Hartmut Surmann, Bernd Möller, and Karl Goser. A distributed selforganizing fuzzy rule-based system. In Neural Networks & their Applications: Proceedings of Neuro-Nimes ’92, pages 187–194, 1992.

Surmann:1992:AFC

[STUG92] Hartmut Surmann, Thomas Tauber, Ansgar Ungering, and Karl Goser. Architecture of a fuzzy-controller based on field programmable gate arrays. In 2nd International Workshop on Field-Programmable Logic and Applications, Vienna, 31 August–2 September 1992.

Surmann:1995:FRB

[SU95] Hartmut Surmann and Ansgar P.

Ungering. Fuzzy-rule-based systems on general purpose processors. IEEE MICRO, Special issue on fuzzy systems, pages 40–48, August 1995.

Surmann:1992:FCV

[SUG92] Hartmut Surmann, Ansgar Ungering, and Karl Goser. Fuzzy-Controller mit VLSI-Pipeline Architektur für hohe Datenraten. In ITG-Fachbericht 119: Mikroelektronik für die Informationstechnik, Stuttgart, pages 195–200, 4– 6 March 1992.

Surmann:1993:OFC

[SUG93] Hartmut Surmann, Ansgar Ungering, and Karl Goser. Optimized fuzzy controller architecture for field programmable gate arrays. In Field-Programmable Gate Arrays: Architectures and tools for Rapid Prototyping, number 705 in Lecture Notes in Computer Science, pages 124–133. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1993. URL http:

//borneo.gmd.de/EIA/papers/
robot_papers.html.

Surmann:1994:WKH

[SUKG94] Hartmut Surmann, Ansgar P. Ungering, Thorsten Kettner, and Karl Goser. What kind of hardware is necessary for a fuzzy rule based system. In FUZZ-IEEE World Congress on Computional Intelligence, Orlando, pages 274–278, 26 June–2 July

1994. URL http://borneo.gmd.
de/EIA/papers/robot_papers.
html.

Surmann:1989:UDH

[Sur89] Hartmut Surmann. Untersuchung der Hardware-Fehlertoleranz ausgewählter assoziativer Speicherkonzepte basierend auf neuronalen Netzen. Diplomarbeit am lehrstuhl für bauelemente der elektrotechnik, Fakultät für Elektrotechnik, Universität Dortmund, May 1989.

Surmann:1992:SLF

[Sur92] Hartmut Surmann. Schnell-Ladeverfahren für NiCd-Akkumulatoren auf der Basis eines Fuzzy-Controllers. Universität Dortmund, Fakultät für Elektrotechnik, 1 July 1992.

Surmann:1994:VCC

[Sur94] Hartmut Surmann. Verfahren zum Glänzen von Cobalt-Chrom-Dentalgußlegierungen sowie ein Glänzgerät zur Durchführung des Verfahrens. Patent Offenlegungsschrift DE 44 01 896 A1, Bundesrepublik Deutschland, January 1994.

Surmann:1995:AEF

[Sur95] Hartmut Surmann. Automatischer Entwurf von Fuzzy Systemen. Dissertation, Universität Dortmund, Fakultät für Elektrotechnik, VDI-Verlag Düsseldorf, 1995. URL http:

//borneo.gmd.de/EIA/papers/
robot_papers.html.

Surmann:2001:ACF

[Sur01] Hartmut Surmann. About the combination of functional approaches and fuzzy reasoning. In Computational intelligence. Theory and applications. International conference, 7th Fuzzy Days, Dortmund, Germany, October 1–3, 2001. Proceedings, pages 791–799. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 2001. ISBN 3-540-42732-5.

Ungering:1992:AFR

[URSG92] Ansgar Ungering, Peter Roer, Hartmut Surmann, and Karl Goser. Architekturkonzept eines Fuzzy-RISC-Prozessors mit optimiertem Speicherbedarf. In ITG-Fachtagung Re”-ch”-ner”- ge”-stütz”-ter Entwurf und Architektur mikroelektronischer Systeme, Darmstadt, pages 229–237, 23–24 November 1992.